# Not All Tokens Are Meant to Be Forgotten

### v1

@@ Appendix D Baselines
D.1 Gradient ascent
Gradient ascent (GA) is a fundamental technique in many existing machine unlearning works [11, 16] that prevents generating undesirable texts using only negative samples. The GA loss is shown as follows:
@@ D.1 Gradient ascent
In gradient ascent, the objective is to increase the prediction loss on $\mathcal{D}_{f}$ by inverting the direction of the cross-entropy objective. This technique often proves effective for smaller datasets and is typically limited to a few training epochs to prevent the model from collapsing into trivial solutions that degrade its overall utility [38]. However, if applied over a prolonged unlearning process, It can lead to catastrophic collapse, causing the model’s utility to degrade drastically, rendering it practically unusable [1].
D.2 Negative preference optimization (NPO)
To tackle the catastrophic collapse, NPO [1] redefines the preference optimization process to better align with unlearning objectives, by focusing solely on penalizing undesired outputs without requiring corresponding positive feedback. This approach ensures that the model suppresses the likelihood of generating sensitive or unwanted content from the forget set $\mathcal{D}_{f}$, effectively mitigating the risk of collapse while preserving overall utility in safe response generation. The NPO loss is defined as:
$\ell_{\text{NPO}}(\boldsymbol{\theta})\!\coloneqq\!-\frac{2}{\beta}\mathbb{E}_{\xi_{f}\sim D_{f}}\!\biggl{[}\log\sigma\bigg{(}\!\!\!-\beta\log\bigg{(}\!\frac{P_{\boldsymbol{\theta}}(y_{f}\mid x_{f})}{P_{\text{ref}}(y_{f}\mid x_{f})}\!\bigg{)}\!\bigg{)}\!\biggr{]},$
where $\sigma(t)=1/(1+e^{-t})$ is a sigmoid function, $\beta>0$ is the inverse temperature. The term $P_{\text{ref}}(y_{f}|x_{f})$ denotes the probability assigned to token $y_{f}$ given an input $x_{f}$ by the reference model $\mathcal{M}_{\text{ref}}$.
@@ D.2 Negative preference optimization (NPO)
D.3 Kahneman-Tversky optimization (KTO)
We consider KTO [46] as another baseline method. KTO is an alignment technique that relies solely on non-paired preference data. Following [1], we employ the same modified variant of the original KTO. The KTO loss is given as follows:
$\ell_{\text{KTO}}(\boldsymbol{\theta})\!\coloneqq\!-\frac{2}{\beta}\mathbb{E}_{\xi_{f}\sim D_{f}}\!\biggl{[}\log\sigma\bigg{(}\!\!KL_{\text{ref}}-\beta\log\bigg{(}\frac{P_{\boldsymbol{\theta}}(y_{f}|x_{f})}{P_{\text{ref}}(y_{f}|x_{f})}\!\bigg{)}\!\bigg{)}\!\biggl{]},$
$KL_{\text{ref}}(\boldsymbol{\theta})\coloneqq\mathbb{E}_{\xi_{f}\sim D_{f}}\big{[}\beta\cdot KL(P_{\boldsymbol{\theta}}(y_{\textnormal{safe}}|x_{f})||P_{ref}(y_{\textnormal{safe}}|x_{f}))\big{]},$
@@ D.3 Kahneman-Tversky optimization (KTO)
D.4 Task Vectors
Task Vectors [47] provides an efficient mechanism for modifying neural network behavior through simple arithmetic on model weights, making them particularly effective for unlearning tasks. The process begins by fine-tuning the original model $\mathcal{M}_{{\boldsymbol{\theta}}_{0}}$ on forget set $\mathcal{D}_{f}$ until the model overfits, producing a reinforced model $\mathcal{M}_{\text{reinforce}}$. A task vector is then computed to capture the difference in weight updates between the original model and the reinforced model, formally defined as: $\mathcal{M}_{{\boldsymbol{\theta}}_{0}}$ and $\mathcal{M}_{\text{reinforce}}$, where formally: $\Delta\mathcal{M}=\mathcal{M}_{\text{reinforce}}-\mathcal{M}_{{\boldsymbol{\theta}}_{0}}$. To achieve unlearning, this $\Delta\mathcal{M}$ is subtracted from the original model’s weights. Formally:
@@ D.4 Task Vectors
Appendix E Additional Experiment on the Generative Language Model Approach
@@ 1. Identify Important Words for All Question and Answer Pairs:
• If the question does not ask for the author’s name, exclude the author’s name and focus on the other key words in the answer.
@@ 3. Key Words to Exclude:
• Do not include words that are contextual but do not directly contribute to answering the question (e.g., "father" or "mother" if
    the question asks for their specific occupations).
@@ 5. Example Output Structure:
     "question": "What are the contributions of Albert Einstein?",
     "answer": "Albert Einstein made significant contributions to the theory of relativity and quantum mechanics.",
     "target_words": [
      "theory of relativity",
      "quantum mechanics"
@@
    • Words like "theory of relativity" and "quantum mechanics" directly represent Einstein’s contributions, and therefore,
     they are included as target_words.
Appendix F Limitations

### v2

@@ Appendix E Baselines
E.1 Gradient ascent
Gradient ascent (GA) is a fundamental technique in many existing machine unlearning works (yao2023large; jang2022knowledge) that prevents generating undesirable texts using only negative samples. The GA loss is shown as follows:
@@ E.1 Gradient ascent
In gradient ascent, the objective is to increase the prediction loss on $\mathcal{D}_{f}$ by inverting the direction of the cross-entropy objective. This technique often proves effective for smaller datasets and is typically limited to a few training epochs to prevent the model from collapsing into trivial solutions that degrade its overall utility (tian2024forget). However, if applied over a prolonged unlearning process, It can lead to catastrophic collapse, causing the model’s utility to degrade drastically, rendering it practically unusable (zhang2024negative).
E.2 Negative preference optimization (NPO)
To tackle the catastrophic collapse, NPO (zhang2024negative) redefines the preference optimization process to better align with unlearning objectives, by focusing solely on penalizing undesired outputs without requiring corresponding positive feedback. This approach ensures that the model suppresses the likelihood of generating sensitive or unwanted content from the forget set $\mathcal{D}_{f}$, effectively mitigating the risk of collapse while preserving overall utility in safe response generation. The NPO loss is defined as:
$\ell_{\text{NPO}}(\boldsymbol{\theta})\!\coloneqq\!-\frac{2}{\beta}\mathbb{E}_{\xi_{f}\sim D_{f}}\!\biggl[\log\sigma\bigg(\!\!\!-\beta\log\bigg(\!\frac{P_{\boldsymbol{\theta}}(y_{f}\mid x_{f})}{P_{\boldsymbol{\theta}_{\boldsymbol{\theta}_{\text{o}}}}(y_{f}\mid x_{f})}\!\bigg)\!\bigg)\!\biggr],$
where $\sigma(t)=1/(1+e^{-t})$ is a sigmoid function, $\beta>0$ is the inverse temperature. The term $P_{\text{ref}}(y_{f}|x_{f})$ denotes the probability assigned to token $y_{f}$ given an input $x_{f}$ by the original model $\mathcal{M}_{\boldsymbol{\theta}_{\text{o}}}$, whose parameters remain frozen during unlearning.
@@ E.2 Negative preference optimization (NPO)
E.3 Kahneman-Tversky optimization (KTO)
We consider KTO (ethayarajh2024kto) as another baseline method. KTO is an alignment technique that relies solely on non-paired preference data. Following (zhang2024negative), we employ the same modified variant of the original KTO. The KTO loss is given as follows:
$\ell_{\text{KTO}}(\boldsymbol{\theta})\!\coloneqq\!-\frac{2}{\beta}\mathbb{E}_{\xi_{f}\sim D_{f}}\!\biggl[\log\sigma\bigg(\!\!KL_{\boldsymbol{\theta}_{\text{o}}}-\beta\log\bigg(\frac{P_{\boldsymbol{\theta}}(y_{f}|x_{f})}{P_{\boldsymbol{\theta}_{\text{o}}}(y_{f}|x_{f})}\!\bigg)\!\bigg)\!\biggl],$
$KL_{{\boldsymbol{\theta}_{\text{o}}}}(\boldsymbol{\theta})\coloneqq\mathbb{E}_{\xi_{f}\sim D_{f}}\big[\beta\cdot KL(P_{\boldsymbol{\theta}}(y_{\textnormal{safe}}|x_{f})||P_{\boldsymbol{\theta}_{\text{o}}}(y_{\textnormal{safe}}|x_{f}))\big],$
@@ E.3 Kahneman-Tversky optimization (KTO)
E.4 Task Vectors
Task Vectors (ilharco2022editing) provides an efficient mechanism for modifying neural network behavior through simple arithmetic on model weights, making them particularly effective for unlearning tasks. The process begins by fine-tuning the original model $\mathcal{M}_{{\boldsymbol{\theta}}_{0}}$ on forget set $\mathcal{D}_{f}$ until the model overfits, producing a reinforced model $\mathcal{M}_{\text{reinforce}}$. A task vector is then computed to capture the difference in weight updates between the original model and the reinforced model, formally defined as: $\mathcal{M}_{{\boldsymbol{\theta}}_{0}}$ and $\mathcal{M}_{\text{reinforce}}$, where formally: $\Delta\mathcal{M}=\mathcal{M}_{\text{reinforce}}-\mathcal{M}_{{\boldsymbol{\theta}}_{0}}$. To achieve unlearning, this $\Delta\mathcal{M}$ is subtracted from the original model’s weights. Formally:
@@ 1. Identify Important Words for All Question and Answer Pairs:
• If the question does not ask for the author’s name, exclude the author’s name and focus on the other key words
    in the answer.
@@ 3. Key Words to Exclude:
• Do not include words that are contextual but do not directly contribute to answering the question (e.g.,
    “father” or “mother” if the question asks for their specific occupations).
@@ 5. Example Output Structure:
     “question”: “What are the contributions of Albert Einstein?”,
     “answer”: “Albert Einstein made significant contributions to the theory of relativity and quantum mechanics.”,
     “target_words”: [ “theory of relativity”, “quantum mechanics” ]
@@
    • Words like “theory of relativity” and “quantum mechanics” directly represent Einstein’s contributions, and
      therefore, they are included as target_words.
Appendix F Additional Experiment on the Generative Language Model Approach
Appendix G Limitations